Contents

ABSTRACT  E1
1 A first, simple, section heading  E1
2 A funny ANSYS-section with a log  E1
3 Yet another ¿Çäsar Müller não¡  E3
4 Testing Intérnal Jœmps  E3
5 A ßection with ≤  E4
5.0.1 A subsubsection a + b . . . . . . . . . . . . . . . . . . . . E4
5.0.1.1 A paragraph . . . . . . . . . . . . . . . . . . . . E4
6 Testing External Jumps  E5

Abstract 1 A first, simple, section heading

And some text

2 A funny ANSYS-section with a log

Page 1; See page –E2–

E1 Page 2; this is page 2

E2 A dummy URL End Acrobat Reader Switch to full screen Zoom in ∫ dx

3 Yet another ¿Çäsar Müller não¡

4 Testing Intérnal Jœmps

This is some text and this is a target.

E3 Now lets jump to Target. Let’s try going to Section 4

5 A ßection with ≤

xx xx 5.0.1 A subsubsection a + b xx 5.0.1.1 A paragraph xx

A subparagraph xxx Pictures:

Normal

Scaled 0.75 E4 Width 1in height 0.5in

Rotated 50

Rotated -50

6 Testing External Jumps

1. Jump to an external: The jump target should open test7.pdf on page 2,

2. Jump to an external: The jump page 1 should open test7.pdf on page 1. a PS file to launch an application a PS file to launch (2)

E5